OpenStudy (anonymous):

Quadratic Equation: Give exact and approximate solutions to three decimal points. x^2+4x+4=49

jhonyy9 (jhonyy9):

x*2 +4x -45 =0 -4 (+/-) (16+180)*(1/2) -4 (+/-) 14 x_1,_2 = ----------------------- = ------------ 2 2 x_1 =-18/2 = -9 x_2 =10/2 = 5 x_1 = -9 x_2 = 5